OpenStudy (anonymous):

x 5 -5 -3 -7 y=4(x-7)+6 ? ? ? ? is thiz linear , quadratic, or exponential plzzzz help

OpenStudy (anonymous):

plllllzzzzzzzzzzzzzzzzz help

OpenStudy (amistre64):

umm..... what is the question?

OpenStudy (anonymous):

X 5 -5 -3 -7 y=4(x-7)+6 ? ? ? ? is thiz linear ,quadratic , or exponential

OpenStudy (anonymous):

?????????????????????/

OpenStudy (amistre64):

linear; since the greatest power is 1

OpenStudy (amistre64):

but what do the 5, -5, -3, -7 mean?

OpenStudy (amistre64):

(5,-2) (-5,-42) (-3,-34) (-7, -50)

OpenStudy (anonymous):

so is linear

OpenStudy (amistre64):

yes, is linear

OpenStudy (anonymous):

oh thanks

OpenStudy (anonymous):

x 5 -5 -3 -7 y=5(x-2)(x+3) ? ? ? ? is linear ,quadratic or exponential

OpenStudy (amistre64):

that is quadratic; when we mltiply them together we get an x^2 which defines this as quadratic

OpenStudy (amistre64):

quadratic is a fancy name that translates from latin and it mean "we squared it"

OpenStudy (amistre64):

it just so happens that the reason we study quadratics so much is that everything can factor into either linear equations or quadratics

OpenStudy (anonymous):

wait so wat bout 5 -5 -3 & -7

OpenStudy (amistre64):

when you plug those values into your equatoin what numbers do you get?

OpenStudy (amistre64):

5(5-2)(5+3) = ?? (5,??) 5(-5-2)(-5+3) = ?? (-5,??) 5(-3-2)(-3+3) = ?? (-3,??) 5(-7-2)(-7+3) = ?? (-7,??)

OpenStudy (anonymous):

?????????????????/

OpenStudy (anonymous):

ummmmm

OpenStudy (anonymous):

ohhh now i get it lol thank u very much yur helpful